OpenStudy (anonymous):

If a gas occupies 1733 cm3 at 10.0C at what temp will it occupy 950 cm3? Assume that pressure remain constant.

OpenStudy (aaronq):

PV=nRT if P, n and R are constant, then \[\frac{ V _{1} }{ T _{1} }=\frac{ V _{2} }{ T _{} }\]

OpenStudy (anonymous):

1733/10.0=950/?

OpenStudy (aaronq):

T needs to be in Kelvin

OpenStudy (anonymous):

How do I convert to K

OpenStudy (aaronq):

lol 0 Kelvin = -273.15 celsius

OpenStudy (anonymous):

-2730.5

OpenStudy (anonymous):

-2731.5

OpenStudy (aaronq):

no, 10.0C + 273.15 C = 283.15 K

OpenStudy (anonymous):

1733cm3/283.15K?

OpenStudy (aaronq):

yep

OpenStudy (aaronq):

1733/283.15=950/? ->T2=950*283.15/1733

OpenStudy (anonymous):

155.218

OpenStudy (anonymous):

Kcm3?

OpenStudy (aaronq):

T2=950 cm^3*283.15 K/1733cm^3 the cm^3's will cancel each other out.

OpenStudy (anonymous):

would it 155.217K

OpenStudy (aaronq):

ye

OpenStudy (anonymous):

thank you so much;D

OpenStudy (aaronq):

no problem, dude!

OpenStudy (anonymous):

I'm a dudett lol

OpenStudy (aaronq):

oh mah badd, it says 'MAN' in your name :P

OpenStudy (anonymous):

lol its okay part of my last name lol ;D everyone thinks that Lol

OpenStudy (aaronq):

ohhh ok haha

OpenStudy (anonymous):

thanks for your help don't have a chem book makes it hard to work on

OpenStudy (aaronq):

no problem, really. If you want i can give you a link and you can download a pdf